Circular No. 6436

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

GRO J1655–40 AND GRS 1915+105 B. A. Harmon, C. A. Wilson, M. McCollough, S. N. Zhang, W. S.

Paciesas, and C. R. Robinson report for the BATSE Compton Observatory Science Team: “The two superluminal jet sources GRO J1655–40 and GRS 1915+105 are both being detected strongly in hard x-rays by BATSE. GRO J1655–40 was first detected in late May 1996, but has gradually increased in brightness to an average around 0.6 Crab (20–100 keV) beginning about July 1. Occasional daily averages reach 0.9 Crab. The spectrum is similar to that of previous outbursts, exhibiting a power law shape to 200 keV with photon indices of −2.3 to −2.6. GRS 1915+105 has also become active again in hard x-rays, and is undergoing a moderately bright outburst (∼ 200 mCrab in the band 20–100 keV) which began around July 4. This precedes the increased radio flux by ∼ 5 days (IAUC 6432). Recent increases in hard x-ray flux (around June 21 and July 4 onward) have been accompanied by spectral softening. The spectrum, adequately fit by a power law, has a spectral index varying between −2.4 and −3.4.”

SUPERNOVA 1985T IN ANONYMOUS GALAXY R. H. McNaught, Anglo-Australian Observatory, reports his discovery

of an apparent supernova on a J plate taken by F. G. Watson on 1985 June 15 with the U.K. Schmidt telescope, confirmed on an R/I plate pair taken with the same telescope by M. R. S. Hawkins on 1985 June 26. The R plates show the supernova at the same brightness (mag ∼ 18) on both dates; SN 1985T is slightly fainter in I than R, relative to nearby stars. The position of SN 1985T is α = 19^{h}03^{m}28^{s}.54, δ = −65^{o}36′12′′.4 (equinox 2000.0), which is 22′′ east and 21′′ north of the galaxy’s center. No star appears to mag 21 at this position on U.K. Schmidt J plates taken on 1975 June 14 and 1986 Sept. 3, or on R plates taken in 1995/1996; likewise, SN 1985T does not appear on the ESO R (exposure number 5665) or ESO B (number 545) surveys.

COMET C/1995 O1 (HALE-BOPP) Total visual magnitude estimates: June 24.58 UT, 5.6 (D. Seargent,

The Entrance, N.S.W., naked eye); July 8.91, 5.6 (J. D. Shanklin, Versailles, France, 7×50 binoculars); 12.15, 5.7 (J. E. Bortle, Stormville, NY, 10×50 binoculars); 14.28, 5.9 (C. S. Morris, Lockwood Valley, CA, 10×50 binoculars); 15.22, 5.5 (R. Keen, Mt. Thorodin, CO, 3×30 binoculars).

1996 July 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green